OpenStudy (anonymous):

Help please! Can't figure this out and I'm about to punch my laptop screen. Simplify the following expression. (2x + 7) + (-4x2 + 6x + 3) A. 4x + 10 B. -4x2 + 8x + 10 C. 4x2 + 8x + 4 D. -4x2 + 4x + 10

OpenStudy (tkhunny):

Just add the LIKE terms. Like Terms are 2x and 6x Like terms are 7 and 3 -4x^2 has no friends.

OpenStudy (anonymous):

B or C

OpenStudy (tkhunny):

There is no OR, there is only Prove It. Quit guessing. Rule out everything lacking "8x" - That's B and C remaining. Rule out everything lacking a 10 - There goes C. Think through it in a logical fashion. Don't just throw it in the air and hope it lands in the right place. Prove it!

OpenStudy (kl0723):

@lee_actress the best you can do is learn how to do the exercises, this is what we try to do here, because all of this practice will be needed in higher levels of math... do yourself a favor and give it a try ;)

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

hmm... so all were doing is adding the like terms?

OpenStudy (tkhunny):

(2x + 7) + (-4x2 + 6x + 3) I don't see anything else in there.

OpenStudy (anonymous):

okay. So we would add 2 + 7 and 6+3 ??

OpenStudy (tkhunny):

No, those are not "like" terms. I listed them, above. Like Terms are 2x and 6x Like terms are 7 and 3 -4x^2 has no friends.

OpenStudy (anonymous):

oh oh okay. So 2 + 6 =8 and then 7 + 3 = 10...

OpenStudy (tkhunny):

Well, 2x + 6x = 8x. Why do you keep throwing out the x?

OpenStudy (tkhunny):

If you are counting dogs, and you have 3 dogs, and a friend brings 5 more dogs, how many dogs are there?

OpenStudy (anonymous):

Right sorry, 2x + 6x = 8x and 7x + 3x = 10 x ( 8)

OpenStudy (tkhunny):

Why did the x suddenly appear on the 7 and 3? Don't do that. Take them like they are. 3 dogs + 5 dogs = 8 dogs <== Correct 3 dogs + 5 dogs = 8 <== Incorrect 3 + 5 = 8 <== Correct 3 + 5 = 8 dogs <== Incorrect Use the units or type of each set. Don't mix them up.

OpenStudy (anonymous):

Okay I get it :)

OpenStudy (tkhunny):

Woof! :-) Always count dogs. Of course, this will get more difficult when you get to negative numbers and fractional values. Dogs don't loke it when you do that to them.

OpenStudy (tkhunny):

(2x + 7) + (-4x2 + 6x + 3) 2x + 6x + 7 + 3 - 4x^2 8x + 10 - 4x^2 -- Done.

OpenStudy (anonymous):

Oookay.... I'm pretty sure i am understanding this

OpenStudy (anonymous):

so im still stuck on the correct answer though which i know I Shouldnt be

OpenStudy (tkhunny):

Find the one with ALL the same pieces. A. 4x + 10 B. -4x2 + 8x + 10 C. 4x2 + 8x + 4 D. -4x2 + 4x + 10 Okay, A is missing the x^2 piece. Throw out A B. -4x2 + 8x + 10 C. 4x2 + 8x + 4 D. -4x2 + 4x + 10 Now what? You don't have to guess. You can step through it a piece at a time.

OpenStudy (anonymous):

C is the answer! @tkhunny

OpenStudy (anonymous):

Wait the one you wrote wasn't negative

OpenStudy (anonymous):

or was it? Becuase if it was it would be B

OpenStudy (tkhunny):

Excellent. I just copied your original post. C does not have NEGATIVE4x^2. Throw it out!! B. -4x2 + 8x + 10 D. -4x2 + 4x + 10 Now what?

OpenStudy (anonymous):

B !!

OpenStudy (anonymous):

right?